If z = 1/(1-I)(2-3i). Then |z| = Pls answer this From complex number

Grade:11

1 Answers

moumi roy
7 years ago
rationalize the frm...(1+i)(2+3i)/(2-3i)(1-i)(1+i)(2+3i)=-1+5i/(1+1)(4+9)=(-1/2*13)+(5/2*13)..this the complex form.
taking its modulus mod(z)=([-1/26]^2+[5/26]^2)^(1/2)=(1/26)^1/2...this is hopefully the answer,